OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f−1(3)?

OpenStudy (anonymous):

I'm not going to be able stay online for long, but it would be nice to be able to wake up tomorrow and actually understand problems like this! Please help, even if I am not online. Happy Presidents Day. :)

OpenStudy (anonymous):

2^f(x) = x + 4 2^f(x) -4 = x => 2^x - 4 = f-1(x) f-1(3) = 2^3-4 = 8 - 4 = 4. The idea is to write x as a function of y ( or f(x) in this case )